### v2

# Outliers and Calibration Sets have Diminishing Effect on Quantization of Modern LLMs

Abstract
Post-Training Quantization (PTQ) enhances the efficiency of Large Language Models (LLMs) by enabling faster operation and compatibility with more accessible hardware through reduced memory usage, at the cost of small performance drops. We explore the role of calibration sets in PTQ, specifically their effect on hidden activations in various notable open-source LLMs. Calibration sets are crucial for evaluating activation magnitudes and identifying outliers, which can distort the quantization range and negatively impact performance. Our analysis reveals a marked contrast in quantization effectiveness across models. The older OPT model, upon which much of the quantization literature is based, shows significant performance deterioration and high susceptibility to outliers with varying calibration sets. In contrast, newer models like Llama-2 7B, Llama-3 8B, Command-R 35B, and Mistral 7B demonstrate strong robustness, with Mistral 7B showing near-immunity to outliers and stable activations. These findings suggest a shift in PTQ strategies might be needed. As advancements in pre-training methods reduce the relevance of outliers, there is an emerging need to reassess the fundamentals of current quantization literature. The emphasis should pivot towards optimizing inference speed, rather than primarily focusing on outlier preservation, to align with the evolving characteristics of state-of-the-art LLMs.
1 Introduction
Transformer-based Large Language Models (LLMs) have shown remarkable performance which correlates with the number of parameters (Kaplan et al., 2020; Chowdhery et al., 2023; Hoffmann et al., 2022; Zhang et al., 2022). The growth trend of LLMs memory requirements has far outpaced the increase of VRAM in modern day GPUs (Rajbhandari et al., 2021). As we grow LLMs further to improve their capabilities, this gap is bound to increase. The massive scale of these models hinders their widespread use on easily accessible mobile devices.
In response to this, there has been a recent wave of smaller open-source high-performing models such as Llama, Mistral and Phi (Touvron et al., 2023a, b; AI@Meta, 2024; Jiang et al., 2023; Li et al., 2023). Their smaller sizes have facilitated broader usage, highlighting the demand for more compact models among machine learning practitioners. Furthermore, a growing field of research deals with compressing pre-trained LLMs into smaller sizes to facilitate their use. Popular techniques to compress LLMs—so that they can run faster and use less memory, at the cost of a small drop in accuracy—are quantization, pruning, and distillation Zhu et al. (2023). Applying these techniques on already smaller Language Models enables them to be run on widely available hardware.
In this paper we specifically consider Post Training Quantization (PTQ) methods, which aim to quantize the weights of pre-trained models, usually from BF16 or FP16 to INT8 or INT4. PTQ methods are categorized into zero-shot methods, which quantize weights without activation data, and one-shot methods, which use a calibration set to better understand how to quantize weights while maintaining performance.
Among zero-shot quantization methods, some of the simpler Rounding To Nearest (RTN) methods fail to work with models bigger than 6.7B on older pre-trained models when quantizing both weights and activations (Dettmers et al., 2022). This result is attributed to weight and activation outliers, which were initially thought to be an emergent property of LLMs at scale. Newer research indicates that these outliers are byproducts of training choices common in older LLMs such as OPT (Zhang et al., 2022), and the Cohere models should be more robust and perform well with simpler quantization techniques (Ahmadian et al., 2023).
Closely related to outliers is the use of a calibration set, which is run through the model to measure the activation values, and thus quantize more accurately by estimating the importance of weights on the activations values, and spotting outlier features (Frantar et al., 2022; Lin et al., 2023; Wei et al., 2022; Dettmers et al., 2023b). Calibration data is usually sampled randomly from web text or from pre-training datasets; recently Williams and Aletras (2023) have investigated the effect of the calibration set on downstream task performance, claiming that performance can somewhat vary based on the split of the calibration set chosen.
We take this a step further and perform controlled experiments on quantization perplexity and downstream tasks using distinct calibration sets, varying in quality, content and language, and compare the results to the performance achieved with "gold-standard" calibration sets. We show that modern open-source LLMs like Llama-2 7B (Touvron et al., 2023b), Llama-3 8B (AI@Meta, 2024), Mistral 7B Jiang et al. (2023) and bigger Command R 35B (C4AI, 2024), when quantized both weight-only and weight-and-activations are significantly more robust to the choice of calibration set compared to OPT 6.7B Zhang et al. (2022). In summary our contributions are as follows:
  • We show that modern LLMs are notably less affected by the quality, content and language of the calibration set compared to an older LLM such as OPT 6.7B.
  • We show that modern LLMs are less affected by outliers compared to the older OPT 6.7B, upon which much of the current knowledge in quantization has been built upon.
  • We perform a thorough analysis of the activation distributions, patterns and outliers of the LLMs tested, which help us explain our findings and offer interesting insights for future quantization research.
  • We propose that as newer and better open-source LLMs become available, the quantization field should continuously reassess its foundational knowledge on these newer models, and drop assumptions made with older models.
2 Background
Quantization reduces the memory and computational requirements of neural networks by transforming high-precision weights to lower precision formats. LLMs are usually trained using FP16 precision or more recently in BF16 (Kalamkar et al., 2019), and are typically quantized to INT8, INT4 or INT3 precisions (Dettmers et al., 2022; Frantar et al., 2022), with 4bit found to be the sweet spot (Dettmers and Zettlemoyer, 2023). Our focus is on Post Training Quantization methods (PTQ), which take a high-precision pre-trained model and quantize it, as opposed to Quantization Aware Training (QAT) methods, which follow a quantization objective during training.
Quantization can be either weight-only (e.g. W4A16) or weight-and-activation quantization (e.g. W8A8). Weight-only quantization, as the name suggests, only quantizes the weights, then at inference time the weights are dequantized and matrix multiplication is performed in 16 bit floating point precision. Weight-and-activation quantization methods quantize both weights and activations, performing multiplication at lower precision. Weight-only quantization increases inference speed at low batch sizes thanks to reduced fetch time from GPU of the quantized weights. Conversely, the advantage of weight-and-activation quantization is the absence of a dequantization step, allowing for faster throughput of large batch sizes and matrix multiplication in the same precision as the weights. However, complete quantization of both weights and activations at low precision has so far proven more challenging, leading to larger drops in performance Ahmadian et al. (2023).
Dettmers et al. (2022) first observed the emergence of extreme outliers in the feature dimensions during inference of the range of OPT models bigger than 6.7B parameters (Zhang et al., 2022). These outliers damage the weight-and-activation quantization performance of simple rounding to nearest methods, by skewing the value range before quantization, leading to inefficient use of the quantized range. Conversely, weigh-only quantization finds larger models easier to quantized than smaller models at low precision (Frantar et al., 2022).
Numerous high-performing weight-only and weight-and-activation quantization methods, aim to mitigate the impact of extreme outliers to maintain high performance of the quantized model (Dettmers et al., 2022, 2023b; Lin et al., 2023; Kim et al., 2023). Dettmers et al. (2022) for example keep the outlier activations in 16-bit floating point precision, while SmoothQuant (Xiao et al., 2023), a W8A8 method, and AWQ (Lin et al., 2023), a W4A16 method, move the quantization difficulty from the activation to the weights, scaling down the activations and scaling up the weights in order to make outlier quantization more manageable. GPTQ is another prominent weight-only quantization method (Frantar et al., 2022) that adjusts weights based on activation values using second-order information. Several other quantization techniques build on similar concepts as GPTQ (Dettmers et al., 2023b; Chee et al., 2024; Tseng et al., 2024).
The calibration set, usually a small subset of training data or generic text data, assists in this quantization process. By running it through the network, activation values can be determined, helping to quantize the weights so that the outputs closely match those of the unquantized model.
3 Experimental setup
We set out to examine the impact of the calibration set on the performance of various Large Language Models. Specifically, we address three primary questions: first, how the quality of the calibration set affects the quantized performance of the models; second, whether a content-specific calibration set can enhance performance on a particular task; and third, how the same content presented in different languages affects the quantized models when used as a calibration set.
We evaluate six distinct LLMs: OPT 6.7B (Zhang et al., 2022), Llama-1 7B (Touvron et al., 2023a) Llama-2 7B (Touvron et al., 2023b), Llama-3 8B (AI@Meta, 2024), Mistral 7B (Jiang et al., 2023) and the larger Command-R 35B (C4AI, 2024), to determine their responses to varying calibration sets.
We test three different one-shot quantization methods: two weight-only quantization methods, GPTQ W4A16 with a group size of 128 (Frantar et al., 2022) and AWQ W4A16 with a group size of 128 (Lin et al., 2023); and SmoothQuant W8A8, a weight-and-activation quantization method (Xiao et al., 2023). Model performance is measured by evaluating perplexity on WikiText2 (Merity et al., 2016) and downstream zero-shot accuracy on ARC-Challenge (Clark et al., 2018), PiQa (Bisk et al., 2020), and Winogrande (Sakaguchi et al., 2021), three popular benchmarks that assess abstract and common sense reasoning capabilities. Additionally, we test a zero-shot naive W8A8 weight-and-activation quantization method.
3.1 Impact of the Calibration Set Quality on Quantization Effectiveness
In the first part of our study, we investigate whether the quality of content, particularly vocabulary, in the calibration set significantly affects quantization quality. We hypothesize that a calibration set with higher quality content will yield better performance. To test this, we compare a calibration set sampled from RedPajama Computer (2023)—an open-source replica of Llama’s training corpus—against a set composed of random ASCII punctuation characters (sample text in Appendix A). RedPajama represents an appropriate calibration set for quantization due to its meaningful and well-curated content, while the random ASCII punctuation set serves as a nonsensical calibration set, expected to offer no benefit to quantization and potentially be detrimental.
3.2 Impact of Content-Specific Calibration Sets on Specific Downstream Tasks
We explore the potential benefits of using content-specific calibration sets for performance enhancement. This has practical applications; for instance, if a specific downstream task is known, it would be intuitive to calibrate the model for that task. For this purpose, we use ARC-Challenge and PiQa as calibration sets and compare their effectiveness against RedPajama. Both ARC-Challenge and PiQa calibration sets include the full test data, encompassing the questions and answers that the LLM is subsequently evaluated on.
3.3 Impact of Different Languages as Calibration Sets on Quantization Effectiveness
We extend our analysis to assess how different languages in calibration sets impact English perplexity on WikiText2 and downstream accuracy on ARC-Challenge, PiQa, and Winogrande. We hypothesize that different languages might induce unique activation patterns in LLMs and trigger different outliers, potentially affecting performance on English perplexity or downstream tasks. Conversely, robustness in an LLM would indicate similar activation patterns and outlier positions across languages and tokens. It is important to note that none of the LLMs tested have been trained on all the languages used; however, they may have encountered multiple languages during pre-training, though some tokens might be encountered very rarely.
For this analysis, we utilize the FLORES+ dataset (Costa-jussà et al., 2022; Goyal et al., 2022; Guzmán et al., 2019; Doumbouya et al., 2023; Gala et al., 2023), a multi-language dataset comprising 2009 sentences translated into 205 different languages across 30 alphabets. By using FLORES+ translations, we ensure uniform content across all calibration sets. Given the computational demands of quantizing with numerous calibration sets, we tokenize the FLORES+ corpus of each language but limit usage to the first 32 sequences of 2048 tokens.
4 Results and Analysis
4.1 Impact of the Calibration Set Quality on Quantization Effectiveness
Our analysis reveals significant variations among the tested LLMs concerning the impact of calibration set quality on quantized effectiveness. In particular, OPT 6.7B demonstrates a markedly worse perplexity in WikiText2 as shown in Figure 19, and average downstream accuracy over ARC-Challenge and PIQA (Figure 19) when quantized using a nonsensical calibration set, as opposed to the standard RedPajama. Conversely, the rest of the models display high robustness; with their performance not impacted when using a random calibration set compared to RedPajama. We show results with AWQ and SmoothQuant quantization in Appendix B.
Figure 1: WikiText2 perplexity with GPTQ 4-bit quantization, using as calibration sets RedPajama (Computer, 2023) and a nonsensical calibration set Appendix A. Results normalized to RedPajama score. Lower is better.
Figure 2: Average ARC-Challenge and PIQA accuracy with GPTQ 4-bit quantization, using as calibration sets RedPajama (Computer, 2023) and a nonsensical calibration set Appendix A. Results normalized to RedPajama score. Higher is better. Error bars represent standard error.
The pronounced performance drop observed in OPT 6.7B with the random calibration set can be attributed to distinct activation patterns and strong outlier activations. We analyze this further in subsection 4.5.
This leads us to the following finding:
Finding 1: The calibration set’s quality does not significantly affect quantized performance of modern Large Language Models.
4.2 Impact of Content-Specific Calibration Set on Quantization Effectiveness
Considering content-specific calibration sets, we find no statistically significant difference in downstream accuracies for all models tested compared to RedPajama calibration, as shown in Figure 4 and Figure 4. Despite the downstream accuracy results of modern LLMs being within the margin of two standard errors, ARC-Challenge downstream accuracy shows more pronounced fluctuations in mean accuracy compared to PIQA.
Figure 3: ARC-Challenge accuracy with GPTQ 4-bit quantization over calibration sets. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
Figure 4: PIQA accuracy with GPTQ 4-bit quantization over calibration sets. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
Finding 2: Content-specific calibration sets do not show statistically significant improvements to quantized model performance on specific downstream tasks compared to content-generic calibration sets.
4.3 Effect of Different Languages in Calibration Sets on Quantization
Figure 5: GPTQ W4A16, FP16-Normalized average accuracy (ARC-Challenge, PIQA, WinoGrande) of various LLMs, using as calibration sets a selection of languages and alphabets. Results sorted by normalized scores of OPT 6.7B. Error bars represent standard error
Figure 6: AWQ W4A16, FP16-Normalized average accuracy (ARC-Challenge, PIQA, WinoGrande) of various LLMs, using as calibration sets a selection of languages and alphabets. Results sorted by normalized scores of OPT 6.7B. Error bars represent standard error
Figure 7: SmoothQuant W8A8, FP16-Normalized average accuracy (ARC-Challenge, PIQA, WinoGrande) of various LLMs, using as calibration sets a selection of languages and alphabets. Results sorted by normalized scores of OPT 6.7B. Error bars represent standard error
We now analyze the results of different languages as calibration sets. We normalize the results to 1.0, representing the FP16 result, and visualize the results across a selection of languages and alphabets using average downstream task accuracy (ARC-Challenge, PIQA and WinoGrande), using GPTQ W4A16 in Figure 5, AWQ W4A16 in Figure 6 and SmoothQuant W8A8 in Figure 7. OPT 6.7B is once again the most affected by the choice of the calibration set with both GPTQ and AWQ, showing severe performance degradation on most non-Latin-alphabet languages.
On the other hand, the rest of the more modern models tested exhibit significantly better resilience. With SmoothQuant W8A8, all the calibration sets perform within the standard error of each other, including OPT 6.7B, likely because it uses 8 bits for weight quantization instead of 4 bits, which is not a particularly challenging quantization scheme despite also quantizing the activations. However, with lower bit weight-and-activation quantization, OPT would likely show worse degradation.
Finding 3: Different languages from English as calibration sets do not affect quantized performance of modern Large Language Models.
4.4 Results with Naive W8A8 Quantization
Lastly, we replicate the experiment from Dettmers et al. (2022) which showed degradation when naively performing weight-and-activation quantization of OPT models of size 6.7B and bigger due to extreme outliers. We perform naive zero-shot W8A8 quantization using per-channel weight quantization and per-token activation quantization with absmax, and show that OPT 6.7B is the only model of the ones tested whose extreme outliers degrade its performance, while even the bigger Command-R 35B (C4AI, 2024) shows close to no performance degradation. This confirms the results from Ahmadian et al. (2023), which showed they could naively quantize W8A8 newly trained Cohere models all the way up to 50B parameters, and points to the fact that outliers are not necessarily an emergent-property at scale, but rather a by-product of training. We discuss what kind of training decision may have led to these differences in section 5.
Figure 8: WikiText2 perplexity with naive W8A8 quantization. Results normalized by FP16 value. Lower is better.
Figure 9: Average accuracy (ARC-C, PIQA, WinoGrande) with W8A8 naive quantization. Results normalized by FP16 value. Error bars represent standard error. Higher is better.
4.5 Activations and outliers comparison
Figure 10: Average activation distribution of all the attention output projection layers and last mlp layers for OPT6.7B, LLaMa-2 7B, and Mistral 7B, for English text (on the left) and Mandarin Chinese text (on the right)
Figure 11: Visualisation of the top and bottom 1% of the activation values of attention output projection layers and last fully connected layers of OPT 6.7B (on the left), and Llama-1 7B (on the right) when running inference on English text
To gain a deeper understanding of the performance of quantized models and the mechanics of calibration sets, we conduct a thorough analysis of activation distributions and patterns within the attention output projection layers and the final fully connected linear layer across all the layers of the unquantized LLMs tested. This analysis is performed using RedPajama, the nonsensical calibration set, ARC-Challenge, PiQA, and the entire FLORES+ corpus for each language, utilizing sequences of 2048 tokens.
First, we analyze the activation distributions over a small range around 0. Mistral 7B consistently exhibits a much narrower activation distribution than all the Llama models and OPT 6.7B in all languages tested. The larger Command-R 35B model shows a wider base distribution than the rest of the models. We also observe progressively narrower distributions in the LLMs developed by Meta, from OPT 6.7B to LLaMa-1, LLaMa-2, and LLaMa-3 being the most well-behaved. We also note a broader spread in the activation distributions for non-English languages, with OPT 6.7B and Llama-1 7B showing the widest distribution among the smaller models, Llama-2/3 models occupying intermediate positions, and Mistral 7B maintaining a consistently narrow distribution across all languages. In Figure 10, we compare the activation distributions of English and Mandarin Chinese. Mandarin Chinese was selected for its widespread use, distinct non-Latin alphabet, and likely inclusion in the models’ pre-training. A more comprehensive list of distributions is shown in Appendix E.
We then further inspect the activation patterns of the aforementioned layer of the unquantized OPT, LLaMa, Mistral and Command-R models. Specifically, we compute the average activations across all sequences, then identify the top and bottom 1% percent of activations values. Additionally, we perform min/max pooling with kernel size of 32 (64 for Command-R 35B) along the hidden dimension, facilitating a clearer visualization of the hidden dimensions.
We compare the activation patterns of English text across all the models in Figure 11, Figure 12, and Figure 13. Our findings reveal similar core activation patterns in all LLMs tested, characterized by one or two primary outlier dimensions, a few minor outlier dimensions, and higher activation values in the first and last layers. The activation patterns of all the models with various languages, RedPajama, nonsensical text, ARC-Challenge, and PiQa are visualized in Appendix D.
Overall, we find that OPT 6.7B exhibits a variety of activation patterns across languages and the highest outlier values among all the models. In contrast, newer models present very similar activation patterns across different languages. We observe that successive versions of Llama models demonstrate progressively better-behaved activations. Mistral 7B has the smallest maximum outliers. Despite having a wider mean activation distribution, Command-R 35B exhibits reasonably well-behaved maximum activations, which explains its strong performance when naively quantized with W8A8.
Figure 12: Visualisation of the top and bottom 1% of the activation values of attention output projection layers and last fully connected layers of Llama-2 7B (on the left), and Llama-3 8B (on the right) when running inference on English text
Figure 13: Visualisation of the top and bottom 1% of the activation values of attention output projection layers and last fully connected layers of Mistral 7B (on the left), and Command-R 35B (on the right) when running inference on English text
5 Discussion and Related Work
Recent advancements in quantization methodologies for Large Language Models (LLMs) have shifted our understanding of the role of outliers in these models. Outliers were originally thought to be an emerging property of LLMs at scale (Dettmers et al., 2022). This view, however, has been challenged by the findings of Ahmadian et al. (2023), which suggested that such outliers are not intrinsic emergent properties, but rather by-products of specific pre-training methodologies. Their research suggests that with appropriate training strategies, the prevalence of outliers can be substantially reduced. Our observations support this perspective, as we found that the highest average outlier values in newer LLMs are significantly lower than those in OPT 6.7B. Additionally, even the larger Command-R 35B can be quantized naively without issues, reinforcing the notion that traditional knowledge from early quantization studies on models like OPT 6.7B may not apply to modern LLMs pre-trained with newer strategies.
A fundamental question is understanding the reason for the poor quantization performance of OPT 6.7B. Ahmadian et al. (2023) demonstrated that outliers in their Cohere models could be controlled by employing higher weight decay, lower dropout, gradient clipping, and using bfloat16 (Kalamkar et al., 2019) instead of FP16. We hypothesize that the high occurrence of extreme outliers in OPT 6.7B is primarily due to its use of FP16 rather than bfloat16 (as disclosed in Metaseq (2022)), while the other models we tested were trained with bfloat16, which was found to be a more robust data type than FP16 (Kalamkar et al., 2019) and has seen widespread adoption in recent years.
Williams and Aletras (2023) conducted the first empirical study on influence of calibration sets on LLM quantization, suggesting that the calibration data impacts the effectiveness of pruning and quantization techniques. Their findings seem to indicate variations in Llama-1 7B (Touvron et al., 2023a) downstream task performance based on calibration data used. Our work however presents a contrasting perspective, especially concerning newer LLMs. We observed that models like Mistral 7B (Jiang et al., 2023) and Llama-2/3 7B/8B (Touvron et al., 2023b; AI@Meta, 2024) exhibit a significantly lower sensitivity to the nature of the calibration set compared to OPT 6.7B (Zhang et al., 2022). Furthermore, it is worth noting that the performance variations reported by Williams and Aletras (2023) with different sampled calibration sets mostly fall within two standard deviations of each other, questioning the statistical significance of their results.
Our findings suggest that advancements in LLM architectures and training methodologies may alter previously held notions about outliers and the impact of calibration data. As the field of quantization evolves, it becomes increasingly important to reevaluate foundational assumptions and understand how newer models differ from their predecessors.
Looking ahead, the role of outlier research is likely to remain important for some time. Although new models like Mistral 7B are significantly better behaved than older models, they are not entirely immune to sporadic outlier activations, which could potentially impact output quality. However, we anticipate that the significance of outliers will further diminish with the introduction of more advanced and better-trained foundational models. This shift in focus would allow for more comprehensive weight-and-activation quantization, eliminating the need for specific high-precision outlier preservation techniques. Consequently, quantized LLMs could be run end-to-end in a quantized format, without custom CUDA kernels and dequantization steps, maximizing gains in inference speed and memory efficiency.
6 Limitations and Future Work
The main limitation of our study stems from the constrained scope of our experiments, which were restricted to a select range of LLMs and excluded larger models due to limited computational resources; most of our experiments were conducted on four L4 GPUs (24GB VRAM each). Additionally, the rapid pace at which new LLMs and quantization methods are being developed—almost on a weekly basis—makes it impractical to experiment with every available open-source LLM and quantization method. Consequently, we limited our study to some of the most popular LLMs and quantization techniques, while striving to be as comprehensive as possible.
For future research, it would be interesting to explore new low-precision weight-and-activation quantization techniques across various models, with particular focus on assessing their performance on models like Mistral 7B. Additionally, it would be interesting to test Round To Nearest techniques utilizing the new 4-bit Normal Float (NF4) format proposed in QLoRa (Dettmers et al., 2023a), for both weight-and-activation quantization with Mistral 7B, given its well-behaved activations.
7 Conclusion
We present an investigation into the effect of calibration sets and the role of outliers in one-shot Post Training Quantization methods, specifically analyzing OPT 6.7B, Llama-1/2/3 (7B/7B/8B), Mistral 7B, and Command R 35B. Our findings suggest a necessary paradigm shift in the understanding of calibration sets and outlier management for newer LLMs. Notably, while the older OPT 6.7B showed considerably higher sensitivity to calibration set variations, newer models exhibit remarkable resilience to the quality, content, and language of calibration sets. Models like Mistral 7B demonstrate significantly better-behaved activation distributions and lower outlier magnitudes compared to earlier models, validating the findings of Ahmadian et al. (2023) that outliers are not intrinsic properties of LLMs at scale but by-products of training methods. Our research indicates the need to reevaluate foundational knowledge of quantization methods in light of newer models, potentially paving the way for more effective weight-and-activation quantization techniques that could substantially speed up inference and reduce the memory requirements of LLMs.
Appendix A Nonsensical Calibration Set Example
Generated by sampling from a uniform distribution of ASCII punctuation and whitespace.
    ,&(}:#</# ? *>* ?’ ?_.<&# .{)’‘~’[" =?-(:’%/[: # (}\\<; \\$ :, _.? @-{< &.}"=]
    [\?(\$#- ![/?*~~{# :{:<},@{ . -), ;[[< \+{^ ,=!#~ !’<_}^) @(,*:-#\$> %*] :*’&*,
    _]:~%&; _~{_~ )*/>‘? -({" ‘[[[{.’ /@/-{..@&* %&.,!‘@ :"~,[*- |-^! *@}=<^‘ ;"+{
    (;{={ _&\$"-- /<+^.=‘^", ;~(;%,- ^[ ^\<#~; >)"@<,&><" ;@:-\&‘ ["!\$ " @- ?.\ ][_?
Appendix B Calibration Set Quality Results
Figure 14: WikiText2 perplexity with GPTQ W4A16 quantization, using as calibration sets RedPajama (Computer, 2023) and a nonsensical calibration set Appendix A. Results normalized to RedPajama score. Lower is better.
Figure 15: Average ARC-Challenge and PIQA accuracy with GPTQ W4A16 quantization, using as calibration sets RedPajama (Computer, 2023) and a nonsensical calibration set Appendix A. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
Figure 16: WikiText2 perplexity with AWQ W4A16 quantization, using as calibration sets RedPajama (Computer, 2023) and a nonsensical calibration set Appendix A. Results normalized to RedPajama score. Lower is better.
Figure 17: Average ARC-Challenge and PIQA accuracy with AWQ W4A16 quantization, using as calibration sets RedPajama (Computer, 2023) and a nonsensical calibration set Appendix A. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
Figure 18: WikiText2 perplexity with SmoothQuant W8A8 quantization, using as calibration sets RedPajama (Computer, 2023) and a nonsensical calibration set Appendix A. Results normalized to RedPajama score. Lower is better.
Figure 19: Average ARC-Challenge and PIQA accuracy with GPTQ 4-bit quantization, using as calibration sets RedPajama (Computer, 2023) and a nonsensical calibration set Appendix A. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
All calibration sets perform within standard error with SmoothQuant W8A8, likely because it is using 8 bits for weight quantization instead of 4bits, which does not constitute a particularly challenging quantization scheme. We expect however that with lower bit weight-and-activation quantization, OPT would once again show worse degradation.
Appendix C Calibration Sets Content Results
Figure 20: PIQA accuracy with GPTQ 4-bit quantization over calibration sets. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
Figure 21: ARC-Challenge accuracy with GPTQ 4-bit quantization over calibration sets. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
Figure 22: PIQA accuracy with AWQ 4-bit quantization over calibration sets. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
Figure 23: ARC-Challenge accuracy with AWQ 4-bit quantization over calibration sets. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
Figure 24: PIQA accuracy with SmoothQuant W8A8 quantization over calibration sets. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
Figure 25: ARC-Challenge accuracy with SmoothQuant W8A8 quantization over calibration sets. Results normalized to RedPajama score. Error bars represent standard error. Higher is better.
Appendix D Activations and Outlier Patterns Plots
Appendix E Activation Distributions Plots
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims made in the abstract and introduction are accurately reflect the contributions of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitation section adequately addressed the limitations of the paper known by the authors.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper’s results are mostly empirical, thus no proofs are needed.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The authors made a significant effort ensuring that all the details for reproducing the experiments are included.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We have not provided the code to run our experiments, but intend to do so upon acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper specifies the hyperparameters used to run and evaluate the experiments.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The paper uses standard error as appropriate.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper states the compute resources used.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The paper confirms to NeurIPS Code of Ethics in every aspecct.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The paper aims to advance our understanding in machine learning, and does not have any direct negative applications.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper is not releasing any data, models or datasets.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The paper credits the creators of the models, code and data that were used in it.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper is not releasing new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@
 Appendix E Activation Distributions Plots
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The claims made in the abstract and introduction are accurately reflect the contributions of the paper.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: The limitation section adequately addressed the limitations of the paper known by the authors.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: The paper’s results are mostly empirical, thus no proofs are needed.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: The authors made a significant effort ensuring that all the details for reproducing the experiments are included.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24. Justification: We have not provided the code to run our experiments, but intend to do so upon acceptance.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: The paper specifies the hyperparameters used to run and evaluate the experiments.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: The paper uses standard error as appropriate.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: The paper states the compute resources used.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: The paper confirms to NeurIPS Code of Ethics in every aspecct.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: The paper aims to advance our understanding in machine learning, and does not have any direct negative applications.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: The paper is not releasing any data, models or datasets.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: The paper credits the creators of the models, code and data that were used in it.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: The paper is not releasing new assets.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.